# On the foundations and applications of Lorentz-Finsler Geometry

### v1

@@ 2.2.1. From vector spaces to manifolds
Essentially, a Finsler metric, a wind Finsler structure, a cone structure and a Lorentz-Finsler metric on a (connected) manifold $M$ are smooth poitwise assignations (to each point $p\in M$) of, respectively, a Minkowski norm, a wind Minkowski norm, a cone and a Lorentz-Minkowski norm, as defined above. This smooth dependence can be tidely expressed in the case of a Finsler metric $F$ with the following definition. Notice that the latter is also is adapted to our preferences of 2-homogeneity ($L=F^{2}$) and smooth domain, thus using the slit tangent bundle, obtined by removing the 0 section from $TM$.
@@
In case of wind and cone structures, however, it is natural to regard them as hypersurfaces $\Sigma$, $\mathcal{C}$ of $TM$ such that their restrictions $\Sigma_{p}$, $\mathcal{C}_{p}$ resp., to each $T_{p}M$ are wind Minkowski norms and cones. There is, however, a subtlety here, because the smoothness of $\Sigma$ or $\mathcal{C}$ would not be enough to ensure the smooth dependence of the structure with $p$ (see Prop. 2.12 and Fig. 2 in [39]) and an additional transversality condition is required.
@@
Finally, Lorentz-Finsler metrics can be defined as follows (we refer to the comments to Def. 2.8 as well as [99, §3.2] for optimized hypotheses).
@@
$L:\bar{A}\rightarrow\mathds{R}$ is a Lorentz-Finler metric, and then $(M,L)$ a Finsler spacetime, if $\bar{A}$ is the closure of the (pointwise) interior $A$ of a cone domain $\mathcal{C}$ and the restriction $L_{p}$ of $L$ to $\bar{A}\cap T_{p}M$ is a Lorentz-Minkowski norm for each $p\in M$.
@@
From Remarks 2.4 and 2.10, the definition of gradient is straightforward for any smooth function $f:M\rightarrow\mathds{R}$ in the Finsler case and, in the Lorentz-Finsler one, when $f$ is a temporal function (i.e. $df_{p}$ is a timelike form for each $p\in M$) or just whenever $df_{p}$ is a causal form for all $p$.
@@ 2.2.2. Causality and cone triples
Cone structures $\mathcal{C}$ naturally extend the future cones of the relativistic spacetimes, that is, time-oriented Lorentzian manifolds. Consistently with Def. 2.6, a vector $v\in TM\setminus\{\mathbf{0}\}$ is timelike or lightlike when $v\in A$ or $v\in\mathcal{C}=\partial A$; in both cases, they are causal vectors. Consequently, a piecewise smooth curve $\gamma$ is called timelike or causal when so is its velocity everywhere (including the two limits velocities at the breaks). For $p,q\in M$, we write $p\ll q$ (resp. $p<q$) and say that $p$ lies in the chronological (resp strict. causal) past of $q$ or $q$ lies in the chronological (resp strict. causal) future of $p$ if there exists a timelike (resp. causal) $\gamma$ from $p$ to $q$. This permits to define the chronological and causal futures of $p$ as well as its future horismos as, resp.:
@@
(and analogously for the notions on past) [^4] We will use typical notions of relativistic spacetimes as causal spacetime –the one admitting no closed causal curve– or time function $t$ –continuous function strictly increasing on (future-directed) causal curves– for cone structures in the remainder, with no further mention.. This opens the possibility to formally extend all the notions of Causality for spacetimes, including the ladder or hierarchy of spacetimes (see for example, [127]). Moreover:
@@
    $T$ is $\mathcal{C}$-timelike [^6] This hypothesis can be dropped by suitably replacing the Finsler metric $F$ in the item (3) by a wind Finsler structure, see around [39, Fig. 4]., and
@@
For the following result, and discussion, see [99, §2.4, §5]
@@
    (d) $\mathcal{G}$ is smoothable in any arbitrarily small neighborhood of Span(T)(T) maintaining the Lorentz signature of $g$, thus, the smoothening is a Lorentz-Finsler metric defined on the whole slit tangent bundle (in agreement with Beem’s definition).
@@
As we will see, cone triples are especially useful when $\Omega=d\tau$ for a temporal function (recall Remark 2.14) and $T$ is regarded as a physical field of observers.
@@
It is worth pointing out that, given $\mathcal{C}$ and the timelike 1-form $\Omega$, there is a unique timelike direction which crosses the barycenter of $\bar{A}\cap\Omega^{-1}(t)$ for all $t>0$. However, the above generality for $T$ (even amplified as suggested in footnote 6) will be convenient.
In particular, the smoothoned metric $\mathcal{G}$ obtained from different choices of $T$ in Th. 2.17 (last item) exemplifies the high-non uniqueness of the Lorentz-metrics compatible with a single cone structure.
@@ 2.3.1. Finsler distance and geodesics
Extending Riemannian notions, the length and energy of a curve $\gamma:[a,b]\rightarrow M$ for a Finsler metric $L=F^{2}$ are, respectively,
@@
$\ddot{x}^{k}+\gamma_{ij}^{k}(x,\dot{x})\dot{x}^{i}\dot{x}^{k}=0.$
Here, the formal Christoffel symbols $\gamma^{k}_{ij}$ are computed with the same expression as in the (semi-) Riemannian case, but applied now to the fundamental tensor $g$. However, notice that $g=g_{ij}(x,y)$, depends not only on the point $p\in M$ we are considering (with coordinates $x=(x^{1},\dots x^{n})$), but also on the oriented direction of the tangent vector $v_{p}=\sum_{i}y^{i}\partial_{x^{i}}$ (with natural coordinates $(x,y)$ in $TM$). As in the Riemannian case, Finslerian geodesics locally minimize the distance.
@@ 2.3.2. The exponential map, local and global causality
Using geodesics, one can define the exponential map at a point $\exp_{p}$ for any semi-Finsler metric in a similar way as in the Riemannian case, and the existence of convex neighborhoods is ensured by a result of Whitehead [167]. However, the lack of regularity at 0 of $F$ implies that the symbols $\gamma_{ij}^{k}(x,y)$ may not even be continuously extended to $y=0$ (indeed, $\exp_{p}$ is smooth at 0 only in very particular cases, see [2]). Anyway, the exponential map $\exp$ is always [^7] Even though this might be regarded as common knowledge, an explicit proof does not seem to have been published until H. Koehler’s article [105, Proposition 4.1], where S. Ivanov is cited as the author. $C^{1,1}$ and, following R. Bryant [^8] See the link of MathOverflow (date 23/06/2025): https://mathoverflow.net/questions/479094/how-badly-does-the-geodesic-exponential-map-fail-to-be-c2-on-finsler-manifold., all the derivatives can be also bounded. Indeed, there exists a smooth map $E:(-\epsilon,\epsilon)\times\Sigma_{p}\rightarrow\mathds{R}^{n}$ such that
@@
is a smooth geodesic. Such estimates can be transmited to the Lorentz-Finsler case, see the proof of [95, Lemma 6.6].
Then, the extension of this Finsler setting to the Lorentz-Finsler one turns out similar to the extension from Riemann to Lorentz manifolds, even though it is technically more involved, see [1, 37, 122, 125] (notice that some of these references also sharpen the usual Lorentzian regularity). In particular, when $L$ is Lorentz-Finsler, timelike geodesics become the local maximizers of the energy functional. Morever, lightlike $L$-pregeodesics agree with cone geodesics, thus, the former depends only on the associated cone structure [^9] It is worth pointing out that the conformal invariance of lightlike pregeodesics is valid in any signature (where causality properties cannot be claimed). This occurs both, in the semi-Riemannian and in the semi-Finsler cases (in the latter, even anisotropically conformal changes are included), see [99]. $\mathcal{C}$. The regularity of $\exp$ is enough to ensure that, locally, the geometry and causality of a Lorentz-Finsler manifold mimics the one of a Lorentz Minkowski norm (thus, the one of a Lorentz-Minkowski scalar product) in the sense of, for example, [137, Prop. 5.34, Lemma 14.2]). Then, the basic causality relations $I^{\pm},J^{\pm},E^{\pm}$ (see (4)) and time-separation (Lorenzian distance) maintain the properties for relativistic spacetimes. The causal ladder is also consistently defined and mimicks the standard Lorentzian one.
@@ 2.4. Nonlinear, anisotropic and Finsler connections
In this subsection, our basic reference is [102] (which takes into account progress in previous [90, 91]), apart from standard books on Finsler Geometry such as [8, 153].
@@ 2.4.1. Geodesic spray
Starting at geodesics, a number of geometric structures emerge (we will follow [102, §6.1]).
First, the geodesic equation is naturally associated with the geodesic spray $G$, a vector field defined on the tangent space to the cone domain ($T\bar{A}$ included in the second order tangent space T(TM)T(TM)) [^10] The following notions are defined locally, so that they make sense open subsets of $TM$. This includes not only $A$ but also $\bar{A}$ because $L$ can be locally extended beyond $\bar{A}=\mathcal{C}$ (recall the last item below Def. 2.8).. Using any natural coordinates $x^{i},y^{i}$ for $TM$ (induced from the coordinates $x^{i}$ on $M$), $G$ is defined as:
@@ 2.4.3. Cartan tensor
where $C_{ijk}=\partial_{y^{k}}g_{ij}/2$ are coordinates of the Cartan tensor $C$ of $L$. Notice that in the semi-Riemannian case $g_{ij}$ depends only on the point, thus $C_{ijk}=0$, and, in general, at each point $p$, $C$ is the differential (vertical derivative) of $g_{ij}$ restricted to $T_{p}M$ that is, $C_{p}$ measures how far $g_{ij}(x(p),y)$ is from being semi-Riemannian.
@@ 2.4.4. Anisotropic connections
where $\mathring{\Gamma}_{ij}^{a}(x,y):=\partial_{y^{j}}N^{a}_{i}(x,y)$ is positive 0-homogeneous in $y$, that is, $\mathring{\Gamma}_{ij}^{a}(x,\lambda y)=\mathring{\Gamma}_{ij}^{a}(x,y))$ for $\lambda>0$. One could define a covariant derivative by using the nonlinear connection of a (causal) vector field $Z$ in the direction of a coordinate vector, namely:
@@
Notice, however, that this covariant derivative would not be linear in $Z$ because $\mathring{\Gamma}_{ij}^{a}$ is also dependent of VV. [^11] In [102, §7]), this derivation is used to define the (non-linear) parallel transport $Z$ of a causal vector $Z_{x}$ along a curve $\gamma$ and, then, the linear transport along $\gamma$ using $\nabla^{Z}$ (defined below)..
@@
Then, the functions $\Gamma_{ij}^{a}(x)$ obey the standard rule of transformation of the Christoffel symbols, and the non-linear connection becomes affine (or linear). The expression (12) can be seen as a covariant derivative $\nabla$ on vector fields $(X,Y)\mapsto\nabla_{X}Y$ satisfying the expected elementary properties of linarity and Leibniz rule. In particular, it turns out that $(\nabla_{X}Y)_{p}$ only depends of the value of $X$ at $p$, i.e., $X_{p}=\bar{X}_{p}\Longrightarrow(\nabla_{X}Y)_{p}=(\nabla_{\bar{X}}Y)_{p}$.
An anisotropic connection on the bundle $\bar{A}\rightarrow M$ can be defined as a natural extension of the previous covariant derivative. Now, whenever a (causal) vector field $Z$ is prescribed, one will have an affine connection,
$(X,Y)\rightarrow\nabla^{Z}_{X}Y,$
depending smoothly on $Z$. Its precise definition can be seen in [99, §3], and it implies that $\nabla^{Z}_{X}Y$ depends only of $Z_{p}$. Indeed, $\nabla_{X}Y$ can be formally defined as a vertical vector field on $\bar{A}$ (satisfying the natural formal rules of linearity and Leibniz in $X,Y$) and, thus, identifiable with an assignation
@@
Summing up, $\nabla^{Z}_{X}Y$ becomes a vector field on $M$ which at each point $p$ depends on $Z_{p}$ instead of the whole $Z$.
In coordinates, an anisotropic connection is determined by its Christoffel symbols $\Gamma^{a}_{\,\,ij}(x,y)$ given by
$\nabla^{V}_{\partial_{i}}\partial_{j}=\Gamma^{a}_{\,\,ij}(V)\partial_{a}\;.$ (13)
They depend on the oriented direction but satisfy the usual cocycle rule
$\bar{\Gamma}^{a}_{\,\,ij}(\bar{x}(x,y),\bar{y}(x,y))=\frac{\partial\bar{x}^{a}}{\partial x^{m}}(x)\left(\frac{\partial^{2}x^{m}}{\partial\bar{x}^{i}\partial\bar{x}^{j}}(x)+\frac{\partial x^{k}}{\partial\bar{x}^{i}}(x)\,\frac{\partial x^{l}}{\partial\bar{x}^{j}}(x)\,\Gamma^{m}_{\,\,kl}(x,y)\right)$ (14)
Next, we will see that any non-linear connection $N^{a}_{i}$ has a naturally associated anisotropic connection $\Gamma^{a}_{\,\,ij}$ but, when $N^{a}_{i}$ comes from a semi-Finsler metric, a second anisotropic conection appears.

### v2

@@ 2.2.1. From vector spaces to manifolds
Essentially, a Finsler metric, a wind Finsler structure, a cone structure and a Lorentz-Finsler metric on a (connected) manifold $M$ are smooth poitwise assignations (to each point $p\in M$) of, respectively, a Minkowski norm, a wind Minkowski norm, a cone and a Lorentz-Minkowski norm (defined above). This smooth dependence can be tidely expressed in the case of a Finsler metric $F$ with the following definition. Notice that the latter is also is adapted to our preferences of 2-homogeneity ($L=F^{2}$) and smooth domain, thus using the slit tangent bundle, obtined by removing the 0 section from $TM$.
@@
In case of wind and cone structures, however, it is natural to regard them as hypersurfaces $\Sigma$, $\mathcal{C}$ of $TM$ such that their restrictions $\Sigma_{p}$, $\mathcal{C}_{p}$ to each $T_{p}M$ are wind Minkowski norms and cones, resp. There is, however, a subtlety here, because the smoothness of $\Sigma$ or $\mathcal{C}$ would not be enough to ensure the smooth dependence of the structure with $p$ (see Prop. 2.12 and Fig. 2 in [41]) and an additional transversality condition is required.
@@
Consistently, $\mathcal{C}$ contains the lightlike vectors, the interior (or cone domain) $A$ of $\mathcal{C}$ is the union of the interiors $A_{p}$ at all $p\in M$, that is, $A$ contains all the timelike vectors, $\bar{A}=A\cup\mathcal{C}$ contains the causal vectors, and $A^{*},\bar{A}^{*}$ and $\mathcal{C}^{*}$ contain, resp., the timelike, causal and lightlike forms.
Finally, Lorentz-Finsler metrics can be defined as follows (we refer to the comments to Def. 2.8 as well as [107, §3.2] for optimized hypotheses).
@@
$L:\bar{A}\rightarrow\mathds{R}$ is a Lorentz-Finler metric, and then $(M,L)$ a Finsler spacetime, if $\bar{A}$ is the union of a cone structure $\mathcal{C}$ and its interior $A$ and the restriction $L_{p}$ of $L$ to $\bar{A}\cap T_{p}M$ is a Lorentz-Minkowski norm for each $p\in M$.
@@
From Remarks 2.4 and 2.11, the definition of gradient is straightforward for any smooth function $f:M\rightarrow\mathds{R}$ in the Finsler case and, in the Lorentz-Finsler one, when $f$ is a temporal function (i.e. $df_{p}$ is a timelike form for each $p\in M$) as well as when $df_{p}$ is a causal form for all $p$.
@@ 2.2.2. Causality and cone triples
Cone structures $\mathcal{C}$ naturally extend the future cones of the relativistic spacetimes, that is, time-oriented Lorentzian manifolds. Accordingly, piecewise smooth curves $\gamma$ are called timelike or causal when so is its velocity everywhere (including the two limits velocities at the breaks). For $p,q\in M$, we write $p\ll q$ (resp. $p<q$) and say that $p$ lies in the chronological (resp strict. causal) past of $q$ or $q$ lies in the chronological (resp strict. causal) future of $p$ if there exists a timelike (resp. causal) $\gamma$ from $p$ to $q$. This permits to define the chronological and causal futures of $p$ as well as its future horismos as, resp.:
@@
(and analogously for the notions on past). This opens the possibility to formally extend all the notions of Causality for spacetimes, including the ladder or hierarchy of spacetimes (see for example, [137]). Thus, ee will use typical notions of relativistic spacetimes such as causal spacetime –the one admitting no closed causal curve– or time function $t$ –continuous function strictly increasing on (future-directed) causal curves– for cone structures in the remainder, with no further mention. Moreover:
@@
    $T$ is $\mathcal{C}$-timelike [^4] This hypothesis can be dropped by suitably replacing the Finsler metric $F$ in the item (3) by a wind Finsler structure, see around [41, Fig. 4]., and
@@
For the following result, and discussion, see [107, §2.4, §5]
@@
    (d) $\mathcal{G}$ is smoothable in any arbitrarily small neighborhood of Span(T)(T) maintaining the Lorentz signature of $g$. Thus, the smoothening is a Lorentz-Finsler metric defined on the whole slit tangent bundle (in agreement with Beem’s definition).
@@
As we will see, cone triples are especially useful when $\Omega=d\tau$ for a temporal function (see definition in Remark 2.15) and $T$ is regarded as a physical field of observers.
@@
It is worth pointing out that, given $\mathcal{C}$ and the timelike 1-form $\Omega$, there is a unique timelike direction which crosses the barycenter of $\bar{A}\cap\Omega^{-1}(t)$ for all $t>0$. However, the above generality for $T$ (even amplified as suggested in footnote 4) will be convenient.
In particular, the smoothened metric $\mathcal{G}$ obtained from different choices of $T$ in Th. 2.18 (last item) exemplifies the high-non uniqueness of the Lorentz-Finsler metrics compatible with a single cone structure.
@@ 2.3.1. Finsler distance and geodesics
Extending Riemannian notions, the length and energy functionals on curves $\alpha:[a,b]\rightarrow M$ for a Finsler metric $L=F^{2}$ are, respectively,
@@
$\ddot{x}^{k}+\gamma_{ij}^{k}(x,\dot{x})\dot{x}^{i}\dot{x}^{k}=0.$ (5)
This expression can be derived directly (see for example [52, Section 3]) and will agree with the Lagrange equations for the Lagrangian $L/2$ in §2.5. Here, the formal Christoffel symbols $\gamma^{k}_{ij}$ are computed with the same expression as in the (semi-) Riemannian case, but applied now to the fundamental tensor $g$. However, notice that $g=g_{ij}(x,y)$, depends not only on the point $p\in M$ we are considering (with coordinates $x=(x^{1},\dots x^{n})$), but also on the oriented direction of the tangent vector $v_{p}=\sum_{i}y^{i}\partial_{x^{i}}$ (with natural coordinates $(x,y)$ in $TM$). As in the Riemannian case, Finslerian geodesics locally minimize the distance.
@@ 2.3.2. The exponential map, local and global causality
Using geodesics, one can define the exponential map at a point $\exp_{p}$ for any semi-Finsler metric in a similar way as in the Riemannian case, and the existence of convex neighborhoods is ensured by a result of Whitehead [183]. However, the lack of regularity at 0 of $F$ implies that the symbols $\gamma_{ij}^{k}(x,y)$ may not even be continuously extended to $y=0$ (indeed, $\exp_{p}$ is smooth at 0 only in very particular cases, see [3]). Anyway, the exponential map $\exp$ is always [^5] Even though this might be regarded as common knowledge, an explicit proof does not seem to have been published until H. Koehler’s article [113, Proposition 4.1], where S. Ivanov is cited as the author. $C^{1,1}$ and, following R. Bryant [^6] See the link of MathOverflow (date 23/06/2025): https://mathoverflow.net/questions/479094/how-badly-does-the-geodesic-exponential-map-fail-to-be-c2-on-finsler-manifold., all the derivatives can be also bounded. Indeed, there exists a smooth map $E:(-\epsilon,\epsilon)\times\Sigma_{p}\rightarrow\mathds{R}^{n}$ such that
@@
is a smooth geodesic. Such estimates can be transmited to the Lorentz-Finsler case, see the proof of [100, Lemma 6.6].
Then, the extension of this Finsler setting to the Lorentz-Finsler one turns out similar to the extension from Riemann to Lorentz manifolds, even though it is technically more involved, see [1, 39, 132, 135] (some of them sharpening usual Lorentzian regularity). In particular, when $L$ is Lorentz-Finsler, timelike geodesics become the local maximizers of the energy functional. Morever, lightlike $L$-pregeodesics agree with cone geodesics, thus, the former depends only on the associated cone structure [^7] It is worth pointing out that the conformal invariance of lightlike pregeodesics is valid in any signature (where causality properties cannot be claimed). This occurs both, in the semi-Riemannian and in the semi-Finsler cases (in the latter, even anisotropically conformal changes are included), see [107]. $\mathcal{C}$. The regularity of $\exp$ is enough to ensure that, locally, the geometry and causality of a Lorentz-Finsler manifold mimics the one of a Lorentz Minkowski norm (thus, the one of a Lorentz-Minkowski scalar product) in the sense of, for example, [150, Prop. 5.34, Lemma 14.2]). Then, the basic causality relations $I^{\pm},J^{\pm},E^{\pm}$ (see (4)) and time-separation (Lorenzian distance) maintain the properties for relativistic spacetimes. The causal ladder is also consistently defined and mimicks the standard Lorentzian one.
@@ 2.4. Nonlinear, anisotropic and Finsler connections
Next, our basic reference is [110] (which takes into account progress in previous [98, 99]), apart from standard books on Finsler Geometry such as [10, 167].
For the tangent bundle $TM\rightarrow M$, natural coordinates $(x,y)$ $(x^{1},\dots x^{n},$ $y^{1},\dots y^{n})$, induced from coordinates $x=(x^{1},\dots,x^{n})$ on $M$, are used. The derivatives along the fiber (type $\partial_{y^{k}}$) will be called vertical. We will consider $\bar{A}\subset TM$, the (causal) domain of a Lorentz-Finsler metric $L$, but also applies to the slit tangent bundle $TM\setminus\mathbf{0}$ of a (semi-)Finsler one.
@@ 2.4.1. Geodesic spray
Starting at geodesics, a number of geometric structures emerge by taking vertical derivatives (then, decreasing the order of homogeneity), we will follow [110, §6.1]. The geodesic equation (5) is naturally associated with the geodesic spray $G$, a vector field defined on the tangent space to the cone domain ($T\bar{A}$ included in the second order tangent space T(TM)T(TM)) [^8] The following notions are defined locally, so that they make sense on open subsets of $TM$. This includes not only $A$ but also $\bar{A}$ because $L$ can be locally extended beyond $\bar{A}=\mathcal{C}$ (recall the last item below Def. 2.8).. $G$ is defined as:
@@ 2.4.3. Cartan tensor
where $C_{ijk}=\partial_{y^{k}}g_{ij}/2$ are coordinates of the Cartan tensor $C$ of $L$. Notice that, in the semi-Riemannian case $g_{ij}$ depends only on the point, thus $C_{ijk}=0$. In general, at each point $p$, $C$ is the differential (vertical derivative) of $g_{ij}$ restricted to $T_{p}M$ that is, $C_{p}$ measures how far $g_{ij}(x(p),\cdot)$ is from being semi-Euclidean.
@@ 2.4.4. Anisotropic connections
where $\mathring{\Gamma}_{ij}^{a}(x,y):=\partial_{y^{j}}N^{a}_{i}(x,y)$ is positive 0-homogeneous in $y$, that is, $\mathring{\Gamma}_{ij}^{a}(x,\lambda y)=\mathring{\Gamma}_{ij}^{a}(x,y)$ for $\lambda>0$. One could define a covariant derivative by using the nonlinear connection of a (causal) vector field $Z$ in the direction of a coordinate vector, namely
@@
Notice, however, that this covariant derivative would not be linear in $Z$ because $\mathring{\Gamma}_{ij}^{a}$ is also dependent of y=Z(x)y=Z(x). [^9] Such a derivative has some applications. For example, in [110, §7]), it is used to define a (non-linear) parallel transport $Z$ for a causal vector $Z_{x}$ along a curve $\gamma$, regarding $Z(x)$ as an unknown. Then, the so-obtained parallely tranported $Z$ is used to define a linear parallel transport along $\gamma$ by using $\nabla^{Z}$ defined below.
@@
Then, the functions $\Gamma_{ij}^{a}(x)$ obey the standard rule of transformation of the Christoffel symbols, and the non-linear connection becomes affine (or linear). The expression (13) can be seen as a covariant derivative $\nabla$ on vector fields $(X,Y)\mapsto\nabla_{X}Y$ satisfying the expected elementary properties of linearity and Leibniz rule. In particular, it turns out that $(\nabla_{X}Y)_{p}$ only depends of the value of $X$ at $p$, i.e., $X_{p}=\bar{X}_{p}\Longrightarrow(\nabla_{X}Y)_{p}=(\nabla_{\bar{X}}Y)_{p}$.
An anisotropic connection on the bundle $\bar{A}\rightarrow M$ can be defined in coordinates, by using functions $\Gamma^{a}_{\,\,ij}(x,y)$ which depend on the oriented direction but satisfy the usual cocycle rule
$\bar{\Gamma}^{a}_{\,\,ij}(\bar{x}(x,y),\bar{y}(x,y))=\frac{\partial\bar{x}^{a}}{\partial x^{m}}(x)\left(\frac{\partial^{2}x^{m}}{\partial\bar{x}^{i}\partial\bar{x}^{j}}(x)+\frac{\partial x^{k}}{\partial\bar{x}^{i}}(x)\,\frac{\partial x^{l}}{\partial\bar{x}^{j}}(x)\,\Gamma^{m}_{\,\,kl}(x,y)\right).$ (14)
Indeed, now the anisotropic connection $\nabla$ applied to the coordinate vector fields $\partial_{i},\partial_{j}$ is a vertical vector field on $\bar{A}$ defined as:
$\nabla_{\partial_{i}}\partial_{j}|_{(x,y)}=\Gamma^{a}_{\,\,ij}(x,y)\,\partial_{y^{a}}|_{(x,y)}\in\mathcal{V}(TM),$ (15)
with independence of the chosen natural coordinates $(x,y)$ on $\bar{A}$. By imposing usual linearity and Leibniz rules, one defines $\nabla_{X}Y$ for any vector fields $X,Y$ on $M$ obtaining an anisotropic connection $\nabla$ (in coordinates).
Now, let us rewrite (15) in a fashion closer to affine connections. Choose any vector field $Z:M\rightarrow\bar{A}$ (that is, a causal vector field in the Lorentz-Finsler case or any non-vanishing vector field on $M$ in the Finsler case), and consider then the composition
$\nabla_{X}Y\circ Z:M\rightarrow\mathcal{V}(TM).$
Identifying naturally each vertical vector $\nabla_{X}Y\circ Z|_{p}$ with a vector, denoted $\nabla^{Z}_{X}Y|_{p}$ which belongs to $T_{p}M$ we obtain finally: whenever a (causal) vector field $Z$ is prescribed, one has an affine connection,
$(X,Y)\rightarrow\nabla^{Z}_{X}Y,$ (16)
depending smoothly on $Z$. Notice the notation and identifications:
@@
Indeed, (16) can be used to define an anisotropic connection in an abstract (coordinate free) way, the formal definition can be found in [107, §3]. This abstract definition also recovers the fact (evident from above) that $\nabla^{Z}_{X}Y$ depends, at each point $p$, on the local value of $Y$ around $p$ and only on the values of $Z$ and $X$ at $p$.
Next, our aim is to see first that any non-linear connection $N^{a}_{i}$ has a naturally associated anisotropic connection $\Gamma^{a}_{\,\,ij}$ but, when $N^{a}_{i}$ comes from a semi-Finsler metric, a second anisotropic connection appears.